OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4). 11 19 16 8 help plz

OpenStudy (anonymous):

\[g(4)=(4)^2+3.\]

OpenStudy (johnweldon1993):

\[\large g(x) = x^2 + 3 \] \[\large g(4) = (4)^2 + 3 \] It's basically just substituting in 4 for every 'x' you see in your equation

OpenStudy (anonymous):

Thanks I'm not really good at math

OpenStudy (johnweldon1993):

No problem :)